OpenStudy (anonymous):

graph : f(x) = |x-3| + 5 and -f(x) = |-x+3| - 5

OpenStudy (anonymous):

For question 1 First draw y = x. Shift the graph right 3 units. Reflect the graph where y < 0 against the x-axis and you will get a V-shape graph. Shift the entire graph 5 units up to get f(x) = |x - 3| + 5 For question 2 -f(x) = |-x + 3| - 5 f(x) = 5 - |-x + 3| = 5 - |-(x - 3)| = -|x - 3| + 5 Reflect the graph of f(x) = |x - 3| + 5 downwards to get an inverted V-shape graph. The two graphs should be intersecting at the tips of the V's.